mass in grams of 7•10-4 moles c12 h24

1
Expert's answer
2021-01-11T03:54:04-0500

mass = molar mass × amount

m = (7*10-4 mol) * (12*12 + 24*1 g/mol) = (7*10-4 mol) * (168 g/mol) = 0.1176 g